

********************************************************************
UARS WEEKLY STATUS REPORT - 10 MAR 2000
********************************************************************
***********
SOLSTICE
***********
We continue to gather enough data to meet all the SOLSTICE science
objectives during this period of reduced tape recorder operations.
SOLSTICE operates normally, collecting solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible. Level 2 processing continues to be halted due to bugs in the
OA_SSPP_ATT service. CDHF and CSC personnel are working on the problem.
***********
SUSIM
***********
SUSIM continues to gather a near normal set of daily solar scans
scheduled for periods of predicted telemetry.
***********
ACRIM
***********
ACRIM II STATUS REPORT:
ACRIM II continues normal operations with sensor A. Downloading of missing
days of data for CY 98 & 99 is continuing. Preparations are being made for
reprocessing using an adaptation of ACRIMSAT/ACRIM III algorithms.
***********
PEM
***********
The PEM spectrometers continue to be active under PMON control*. AXIS-2
was activated on February 29 to complete the total AXIS activation. At the
present time, PEM energy deposition is not being generated. Recent data
configuration changes have affected PEM software causing multiple failures.
These failures are being reviewed by the PEM team. In order to correct for new
telemetry requirements, PEM will most likely need to update processing code
with a patch delivery. At this time, updated HEPS calibration parameters will
be submitted. These will influence the in-situ energy deposition predictions,
so new versions of PEM energy deposition Level 3 files will be generated.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
HALOE continues successful operations with all engineering parameters
within nominal operating range. Science data processed products continue
to be of excellent quality.
HALOE was power cycled three times for one orbit each on succeeding days
during the peak of the beta cycle this past week. The UARS FOT was very
helpful during this process.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG is working with the Flight Operations Team to
develop schedules for tape recorder operation to obtain
the maximum amount of northern hemisphere MLS data during
the next period of MLS operation.
Routine processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
The CDHF extends "Best Wishes" to Pat Purcell who is leaving the HALOE
group.
STATUS
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
25 days WIN (5.11A) PROD (L1-L3) (535R-407C-14F=14L)
The CDHF is processing a group of days that are needed by HALOE in support
of the SOLVE Mission. (UARS DAYS 3083 - 3105)
UARS CDHF REHOST:
- Successful processing on Rehost System.
Attitude Extraction, Observatory Log, Sensor Extraction, HALOE, SUSIM.
- Testing WINDII V5.11A on the Rehost System this week.
- Continue to test SOLSTICE V4.62 and HRDI V4.4 on the Rehost System.
DAAC DATA:
- Data created in the month of January 2000 has been previewed for
HALOE, NMC, UKMO, and PEM.
- Data created at MLS_RAC for MLS VERSION 5, UARS DAYS 118-121 has been
previewed and loaded into the DAAC TRANSFER TABLES. DAAC needs to pull
this data.
***********
FOT
***********
REPORTING PERIOD
This report covers 26 February, 2000 (Orbit 46233, GMT Day 057) through 3 March
2000 (Orbit 46337, GMT Day 063).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 060/0020 46278 AXIS On
HALOE 062/2208 46322 Off
HALOE 063/0029 46323 On
The beta angle ranged between the angles of 35.2 degrees to 50.3 degrees for
this report period. The beta angle is now increasing toward a maximum angle of
51 degrees on 04 March 1999 (DOY 064).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.0 to 1.3 degrees C. Additional battery performance data for each
battery on 25 Febraury 2000 and 03 March 2000 are provided in Appendix D.
The clock error ranged from -12.4 to 10.4 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% OUT-OF-LIMITS = 00
% ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% FOT = 012
% HALOE = 02
% AXIS = 01
% HRDI = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
% FOT = 01
Three (3) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
- Transponder A: 00/021 - 2287.496849
- Transponder B: 00/060 - 2287.499207
GROUND SYSTEM OPERATION
One (1) operational support problem occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 00/062 show the observatory data loss to be 565 hours, 44
minutes, 32.601 seconds (an increase of 6 minutes 13.328 seconds since the last
report period).
The increase of 6 minutes 13.328 seconds occurred over one (1) day:
- 6 minutes 13.328 seconds on Day 059 due to data hits.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT continued to work on the safehold algorithms. There are a few
mnemonics that need to be investigated before the final algorithms can be
written. Some of the data is defined in pseudos which are not applicable
onboard. The FOT has requested access to all of CDHF7, so that the equations
for the pseudos can be found and reviewed.
Friday, March 3, 2000 (063) - UARS experienced a small anomaly. Two executive
status messages were observed during realtime. The messages received were
processor control error and processor fell behind. The status buffer gave the
message of PROC RQS EXE with a processor id of 3. This status buffer message
corresponds to report 257. According to the UARS Status Buffer Message
Responses Manual, report 257 is titled processor requested while executing or
waiting. The report states that the OBC has become overloaded and cannot keep
up with its scheduled work. The new request is lost, not queued. The FOT
evaluated this as a benign anomaly and followed the preparated Flight Executive
Response: Check the OBC, take a dump of bank 0 for analysis and clear the
executive Report. Operations were not inhibited and the anomaly is closed.
MEETINGS
No meetings this week.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for March 21, 2000 (DOY 081). This
maneuver will be a Forward to Reverse Flight maneuver (Northern Hemisphere
viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 3 March 2000
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
46237 057/0527 00-067 NBTR NBTR OPS
46248 058/0531 00-069 NBTR NBTR OPS
46266 059/0456 00-070 NBTR NBTR OPS
46267 059/0636 00-071 OBC Clk Adj fr 0.00 to 20.6 msec/day
46275 059/1837 00-072 NBTR NBTR OPS
46278 060/0020 00-073 AXIS 1&2 Power ON
46296 061/0355 00-074 OBC Clk Adj fr 20.6 to 0.00 msec/day
46297 061/0541 00-075 NBTR NBTR OPS
46312 062/0643 00-076 NBTR NBTR OPS
46320 062/1844 00-077 NBTR NBTR OPS
46322 062/2208 00-078 HALOE Power OFF
46323 063/0029 00-079 HALOE Power ON
46323 063/0029 00-080 OBC Clk Adj fr 0.00 to 20.6 msec/day
46328 063/0728 00-077 NBTR NBTR OPS
46335 063/1846 00-081 NBTR NBTR OPS
46337 063/2249 00-082 HRDI HRDI MICRO LOAD
_______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 3 March 2000
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
None to report this period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None to report this period.
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 3 March 2000
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
46267 059/06:42 (28 Feb 00) from 0.00 to 20.6 msec/day
46296 061/03:59 (29 Feb 00) from 20.6 to 0.00 msec/day
46323 063/00:59 (02 Mar 00) from 0.00 to 20.6 msec/day
_______________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 3 March 2000
25 Feb 2000 - GMT Day 056 Beta = 35.2 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 -56.0/-89.6 -3.05/-4.31 +0.0/+0.0 16.0 0.000 0.0
2 +89.6/-5.6 +5.79/+4.83 +29.2/-15.2 25.1 1.028 22.5
3 +89.6/+0.0 +5.15/+3.87 +28.8/-16.4 25.1 1.026 22.2
03 Mar 2000 - GMT Day 063 Beta = 50.3 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 -56.0/-78.4 -3.05/-4.31 +0.0/+0.0 16.0 0.000 0.0
2 +33.6/-11.2 +5.47/+3.87 +26.4/-15.6 25.4 1.050 20.1
3 +33.6/-11.2 +4.51/+2.60 +25.6/-15.2 25.3 1.048 19.9
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---------- ----------
no changes V/T 5
______________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 3 March 2000
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-006 FOT 46337 063/2249 N/A Flight Exec Error
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-004 FOT 46142 050/2136 14170 Generic Late Acquisition
00-006 FOT 46337 063/2249 N/A Flight Exec Error
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
No new reports generated
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- ----------------
99-064 FOT 44135 10/8-9/99 N/A HRDI Cal Lamp Errors
99-071 FOT 44398 299/1212 21512 TPF Not Receiv 512
00-005 FOT see list March 00 14170 Generic Late Acq.
SEVNTFW00/09